# Eaman Jahani

### v4

@@ F.1 Hypotheses and Results
    No rescaling: At the user-level, we find strong and statistically significant evidence of an association between performance and our prompting variables, even after adjusting for Benjamini-Hochberg multiple testing. Token sort ratio, cosine similarity with the previous prompt, and frequency of including the previous prompt were positively associated with performance, while embedding variance and the number of topical transitions negatively correlated with performance. Taken together, these findings suggest that more successful users engaged in more exploitation and wrote prompts that were similar to one another. We also found that longer prompts were associated with higher performance, as shown in the main text.
    At the user-prompt level, we find that when the previous performance was poor, higher exploration (or lower cosine similarity with the previous prompt) was associated with improved performance, although extremely high levels of exploration did not improve performance. In contrast, in cases where previous performance was high, higher exploitation monotonically increased performance. Using token sort ratio with the previous prompt as the measure of exploration also generated similar results. However, when using binary measures of exploration, i.e. topical transition or containing the previous prompt, we found that exploitation is associated with higher performance in the next iteration regardless of the bracket. As these are binary measures of exploration, we could not replicate the non-linearity we observed with continuous measures of exploration in the low-performing group. However, we do find that topical transitions (more exploration) leads to an overall drop in performance and the drop becomes larger for higher performance in the previous attempt. We find a similar pattern when measuring exploitation by whether a prompt contains the previous one. Performance is overall higher if the prompt includes the previous one and the improvement is larger as the previous performance gets higher.
    Z-score rescaling: We find the same results as those described for the unscaled cosine similarity above. The main difference is that the non-linearity between performance and continuous measures of exploitation, i.e. TSR ratio and cosine similarity with previous prompt, observed at the user-prompt level becomes starker for the bottom two brackets of previous performance. For low-performing prompts in the previous attempt, the optimal exploitation level is in the middle and the performance greatly deteriorates for high and low levels of exploitation.
    Percentile rank rescaling: We find the same results as described above, however the non-linearity in exploration/exploitation vs performance in the bottom two brackets of previous performance is now even stronger than those observed with Z-score rescaled cosine similarity as the performance measure.
@@
    No rescaling: We find exactly the same results as those described for the unscaled cosine similarity above.
    Z-score rescaling: We find the same results as those described for the Z-score cosine similarity above. In particular, we again find that the non-linearity between performance and continuous measures of exploitation observed at the user-prompt level is stark for the bottom two brackets of previous performance.
    Percentile rank rescaling: We find the same results as described above for cosine similarity percentile rank rescaling.
@@
To test this hypothesis, we estimated a linear model per each demographic trait as the independent variable, treatment arm fixed effects, and the prompting behaviors outlined below and described in-depth in Section B.B.5 as the outcome variables of interest. To account for multiple testing, we adjusted the p-values for all these models according to the Benjamini-Hochberg procedure.
Results: The results of our analysis are as follows. As a reference, the definitions of these dependent variables are provided in Section B.B.5.
@@
  • Prompt embedding variance: We did not find any significant differences across various demographic traits.
@@
  • Strategic Shifts: We observed statistically significant differences based on age, reported programming and instructional writing frequencies. Younger participants, those with low programming frequency and those with some instructional writing frequency demonstrated decreased topical transitions across their prompts.
@@
  • Successive Prompt Token Sort Ratio: We found significant differences by age, education, programming frequency, and imagery/instructional writing frequencies. Older users, those with post-graduate degrees, those with high programming frequency and writing frequency, both precise instructions and imagery, write prompts that are less similar to each other on average.
@@
  • Successive similarity: We found significant differences by age, gender, education, generative AI outlook, programming skill/frequency, instructional/imagery writing frequency and some occupational skills. On average, older users, males, those with post-graduate degrees, those who reported frequently computer programming, those who reported being strong computer programmers, those with a positive outlook on Generative AI, those who frequently use generative AI, those who report frequently writing instructions or imagery, and those with self-reported occupational skills of critical thinking and social perceptiveness write prompts that are less similar to each other (as measured by cosine similarity) on average.
@@
  • Successive Prompt ‘Contains Previous Prompt’ Dummy: We found significant differences by age, gender, outlook towards generative AI, and imagery writing frequency. Older users and those with neutral outlook toward generative AI are less likely to write prompts that exactly contain the previous. In contrast, males and those with some imagery writing skills are more likely to keep using their previous prompt.
@@
Analysis approach: To test this hypothesis, we first calculated the Z-score of each participant’s realized image observed during our experiment relative to the sampling distribution that we approximated for the prompt that generated that image according to the procedure outlined in Section B.B.3 (see Section C for details on calculating the Z-score). This Z-score quantifies the random variation in observed image quality relative to the prompt’s true underlying quality. Images with higher Z-scores represent instances where the realized image quality exceeded expectations based on the prompt, while lower Z-scores indicate instances where the image quality was randomly lower than expected for a given prompt. By examining the relationship between these Z-scores and subsequent performance and prompting behavior, we can assess the causal impact of this stochasticity on user behavior.
To estimate the relationship between this stochasticity on subsequent prompting behavior and performance, we transformed the Z-score into a trichotomous variable where any Z-score less than -0.45 as “lower-than-expected,” between -0.45 and 0.45 as “expected” and greater than 0.45 as “higher-than-expected.” We then perform two-sample t-tests with comparing the three groups’ relative performance. We also estimate a linear model as a robustness check, regressing participant performance on the trichotomous Z-score variable with treatment arm fixed effects. We also repeat this analysis, treating the Z-score as a continuous variable without the trichotomous transformation. Finally, we also perform this analysis at the user level with the trichotomous variable, where only the Z-score of the observed image of the first prompt is measured, and we test if that affects the average performance of all subsequent user attempts.
Results: The results of our analysis are as follows, with each outcome variable clearly labeled.
@@
    No rescaling: We find statistically significant evidence that increases in the Z-score of the image realized in a previous prompt caused an increase in cosine similarity of the next prompt. When we test differences in the trichotomous variable, the difference between the top bracket and the bottom bracket in the realized Z-score is statistically significant, with the top bracket having a higher performance in the next attempt than the bottom bracket. But we do not find significant differences between the bottom and middle brackets and between the top and middle brackets. When we estimate a linear model, we do not find a statistically significant relationship between the observed Z-score of the prompt and performance in the next iteration. This is due to the non-linearity of the relationship between these variables that exists separately in the negative and positive range of realized Z-score. Additionally, when we estimate the effect at the user-level, we find differences between the top and middle brackets of the first prompt realized Z-score. However, these differences are barely significant ($p=0.048$) and do not account for multiple testing. When fitting a linear model at the user-level, we do not find a a statistically significant relationship between the observed Z-score of the first prompt and average performance in the subsequent attempts.
    Z-score rescaling: The results are very similar to those described for the unscaled cosine similarity above. In particular, we find a statistically significant difference between the top bracket and the bottom bracket, with the top bracket having a higher performance in the next attempt than the bottom bracket. We don’t find statistically significant effects when comparing the middle with top or bottom brackets, or when fitting a linear model. In contrast to the unscaled measure, we do not find any statistically significant effects at the user-level, neither when comparing the brackets on the realized Z-score of the first image nor when fitting a linear model between the realized quality of the first prompt and the performance in subsequent attempts.
    Percentile rank rescaling: We find the exact same results as those explained above for Z-score cosine similarity.
@@
    No rescaling: We find similar evidence that increases in the Z-score of the image realized in a previous prompt caused an increase in cosine similarity of the next prompt. The difference between the top bracket and the bottom/middle brackets in the realized Z-score is statistically significant, with the top bracket having a higher performance in the next attempt than both the bottom and middle brackets. But we do not find significant differences between the bottom and middle brackets. When we estimate a linear model, we do not find a statistically significant relationship between the observed Z-score of the prompt and performance in the next iteration, likely due to the non-linearity that we also observed between performance measured as cosine similarity and realized Z-score of the previous prompt. When we estimate the effect at the user-level, we do not find any statistically significant relationship between realized quality of the first prompt and performance in subsequent attempts, either using the trichotomous variable on Z-score or using a linear model.
    Z-score rescaling: The results are very similar to those described for the unscaled dreamsim score above. In particular, we find a statistically significant difference between the top bracket and both the bottom and middle brackets, with the top bracket having a higher performance in the next attempt. We don’t find statistically significant effects when comparing the middle with the bottom bracket, or when fitting a linear model. Similar to the unscaled measure, we do not find any statistically significant effects at the user-level, neither when using the trichotomous variable nor when fitting a linear model between the realized quality of the first prompt and the performance in subsequent attempts.
    Percentile rank rescaling: We find the exact same results as those explained above for Z-score dreamsim similarity, with the only difference that at the user-attempt level, we observe a statistically significant difference only between the top and bottom brackets. The top bracket in realized quality no longer has a statistically significant improvement in the performance of the next attempt when compared to the middle bracket.
@@
    Prompt Length: e find statistically significant evidence that increases in the Z-score of a realized image causes the subsequent prompt to be longer (either in terms of words or characters). When we test the differences using the trichotomous variable, the difference between the top Z-score bracket and the bottom bracket is statistically significant, with on average longer prompts in the top bracket. However, we do not observe statistically significant differences between either the top or bottom brackets with the middle bracket. When we estimate a linear model at the user-attempt level, we find a similar relationship that is both positive and statistically significant. A unit increase in the Z-score realized quality of an image causes the next prompt to be longer by about 0.3 words on average. Finally, when we estimate the effect at the user-level, we do not find the Z-score of the first prompt to have a statistically significant effect on the length of the subsequent prompts, when using the trichotomous variable or when fitting a linear model.
    Successive Similarity: We find statistically significant evidence that increases in the Z-score of a realized image causes an increase in successive similarity of the prompt. When we test the differences using the trichotomous variable, the difference between the top Z-score bracket (better than expected images) and both the bottom (worse than expected images) and middle (about expected images) brackets is statistically significant, with the top bracket leading to higher similarity with the previous prompt. However, we do not observe a statistically significant difference between the bottom bracket with the middle bracket. When we estimate a linear model, we find a similar relationship that is both positive and statistically significant. An increase in the realized quality (Z-score) of an image causes the next prompt to be more similar to the previous one. When we estimate the effect at the user-level, we do not find statistically significant effects of the first prompt realized quality on average successive similarity, neither when using the trichotomous variable nor when fitting a linear model.
    Successive Prompt Token Sort Ratio: The results are identical to the case of successive cosine similarity described above. We find statistically significant evidence that increases in the Z-score of the image realized in a previous prompt causes an increase in the Token Sort Ratio between successive prompts, using both the trichotomous variable (top vs bottom and middel brackets) and the linear model. At the user-level, we do not find statistically significant effects of the Z-score of the first prompt on average token sort ratio of successive prompts.
    Successive Prompt ‘Contains Previous Prompt’ Dummy: We find statistically significant evidence that increases in the Z-score of a realized image causes an increase in the likelihood that the subsequent prompt contains the previous one. When we test the differences using the trichotomous variable, the difference between the top Z-score bracket and the bottom bracket is statistically significant, with the top bracket more likely to exactly include the previous prompt. However, we do not observe statistically significant differences between either the top or bottom bracket with the middle bracket. When we estimate a linear model at the user-attempt level, we find a similar relationship that is both positive and statistically significant. An increase in the realized quality of an image causes the next prompt more likely to include the previous one. Finally, when we estimate the effect at the user-level, we do not find the Z-score of the first prompt to have a statistically significant effect on the average probability of a prompt containing the preceding prompt, when using the trichotomous variable or when fitting a linear model.
@@
Analysis approach:
To test this hypothesis, we conducted three two-sample t-tests that compare the prompting performance and prompting behavior of the three possible pairs of treatment assignments. All outcome variables are specified in the results below. For robustness, we repeat these analyses with ANOVA to test whether any of the three treatment arms has an effect on the same variables. To account for multiple testing, we adjusted the p-values for tests with the Benjamini-Hochberg procedure.
Results: The results of our analysis are as follows:
@@
    Z-score rescaling: We find statistically significant evidence for differences in performance across treatment arms by comparing participants’ first attempt at recreating the target images, their average across all attempts, and their best attempts (all in terms of taskwide Z-scores, as opposed to task-iteration Z-scores). When we compare first iteration performance, those using DALL-E 3 (verbatim) performed better than both those using DALL-E 3 (Revised) and DALL-E 2, although there was no statistically significant difference between the latter two treatment arms. When we compare participants’ average performance, we find the same results in terms of statistical significance (DALL-E 3 verbatim outperforming DALL-E 2 and DALL-E 3 revised, with no statistically significant difference between DALL-E 2 and DALL-E 3 revised). Finally, when we compare participants’ best attempts pairwise across treatment arms, we find statistically significant differences between all three treatment arms: those using DALL-E 3 (verbatim) were, on average, better than those using DALL-E 3 (Revised), who were, on average, better than those using DALL-E 2. Therefore, those using DALL-E 3 (Verbatim) also performed better than those using DALL-E 2. When we test the relationship between the three treatment arms and these three outcome variables (first, average, and best performance) with ANOVA, all are statistically significant.
    Percentile rank rescaling: We find statistically significant evidence for differences in performance across treatment arms by comparing participants’ first attempt at recreating the target images, their average across all attempts, and their best attempts (all in terms of taskwide percentile ranks). When we compare first iteration performance, those using DALL-E 3 (verbatim) performed better than both those using DALL-E 3 (Revised) and DALL-E 2, although there was no statistically significant difference between the latter two treatment arms. When we compare participants’ average performance, we find the same results in terms of statistical significance (DALL-E 3 verbatim outperforming DALL-E 2 and DALL-E 3 revised, with no statistically significant difference between DALL-E 2 and DALL-E 3 revised). Finally, when we compare participants’ best attempts pairwise across treatment arms, we find statistically significant differences between all three treatment arms: those using DALL-E 3 (verbatim) were, on average, better than those using DALL-E 3 (Revised), who were, on average, better than those using DALL-E 2. Therefore, those using DALL-E 3 (Verbatim) also performed better than those using DALL-E 2. When we test the relationship between the three treatment arms and these three outcome variables (first, average, and best performance) with ANOVA, all are statistically significant.
@@
    Z-score rescaling: We find statistically significant evidence for differences in performance across treatment arms by comparing participants’ first attempt at recreating the target images, their average across all attempts, and their best attempts (all in terms of taskwide Z-scores). When we compare first iteration performance, those using DALL-E 3 (verbatim) performed better than both those using DALL-E 3 (Revised) and DALL-E 2, although there was no statistically significant difference between the latter two treatment arms. When we compare participants’ average performance, we find the same results in terms of statistical significance (DALL-E 3 verbatim outperforming DALL-E 2 and DALL-E 3 revised, with no statistically significant difference between DALL-E 2 and DALL-E 3 revised). Finally, when we compare participants’ best attempts pairwise across treatment arms, we find the same pattern of head-to-head statistical significance results. When we test the relationship between the three treatment arms and these three outcome variables (first, average, and best performance) with ANOVA, all are statistically significant.
    Percentile rank rescaling: We find statistically significant evidence for differences in performance across treatment arms by comparing participants’ first attempt at recreating the target images, their average across all attempts, and their best attempts (all in terms of taskwide percentile ranks). When we compare first iteration performance, those using DALL-E 3 (verbatim) performed better than both those using DALL-E 3 (Revised) and DALL-E 2, although there was no statistically significant difference between the latter two treatment arms. When we compare participants’ average performance, we find the same results in terms of statistical significance (DALL-E 3 verbatim outperforming DALL-E 2 and DALL-E 3 revised, with no statistically significant difference between DALL-E 2 and DALL-E 3 revised). Finally, when we compare participants’ best attempts pairwise across treatment arms, we find statistically significant differences between all three treatment arms: those using DALL-E 3 (verbatim) were, on average, better than those using DALL-E 3 (Revised), who were, on average, better than those using DALL-E 2. Therefore, those using DALL-E 3 (Verbatim) also performed better than those using DALL-E 2. When we test the relationship between the three treatment arms and these three outcome variables (first, average, and best performance) with ANOVA, all are statistically significant.
@@
    Mean prompt Length: We found statistically significant differences in prompt length between treatment arms. Participants using DALL-E 2 used significantly shorter prompts compared to both DALL-E 3 (Revised) and DALL-E 3 (Verbatim) groups. There was no significant difference in prompt length between the two DALL-E 3 groups. ANOVA results confirmed a significant effect of DALL-E version on prompt length.
    Aggregate Similarity: We found no statistically significant differences between any of the treatment arms. This suggests that the overall variability in prompts was similar across all three DALL-E versions. ANOVA results confirmed no significant effect of DALL-E version on this measure.
    Successive Similarity: Analysis of the average cosine similarity between successive prompts revealed no statistically significant differences between the treatment arms. ANOVA results confirmed no significant effect of DALL-E version on successive prompt similarity.
    Successive Prompt Token Sort Ratio: We found no statistically significant differences between the treatment arms. ANOVA results confirmed no significant effect of DALL-E version on this measure.
    Successive Prompt ”Contains Previous Prompt” Dummy: Examining the probability of a current prompt being a superset of the previous prompt showed no statistically significant differences between any of the treatment arms. ANOVA results confirmed no significant effect of DALL-E version on this measure.
@@
Analysis approach:
To test this hypothesis, we conducted two analyses. First, would conducted F-tests comparing the variance of participant performance and prompting behaviors between all 3 pairs of models. Second, we estimated the quantile treatment effects (QTEs) between all 3 pairs models on participant performance and prompting behaviors. We also pre-registered our intent to visually inspect whether the QTEs we observe are consistent with dispersion/“inequality” being reduced or increased when participants use different models (e.g., positive effects for low quantiles and negative/null effects for high quantiles would be consistent with inequality reduction).

### v5

@@ G.1 Hypotheses and Results
    • No rescaling: At the user level, we found strong and statistically significant associations between performance and prompting strategies, even after Benjamini-Hochberg adjustment for multiple testing. Specifically, token sort ratio, cosine similarity with the previous prompt, and frequency of including the previous prompt were all positively associated with performance. Conversely, embedding variance and number of topical transitions showed negative correlations with performance. Together, these findings indicate that more successful users engaged in greater exploitation, writing prompts that exhibited higher similarity to one another. As noted in the main text, we also found that longer prompts were associated with higher performance.
      At the user-prompt level, we observed an interesting pattern: when previous performance was poor, moderate exploration (lower cosine similarity with the previous prompt) was associated with improved subsequent performance, though extremely high levels of exploration did not yield additional benefits. In contrast, when previous performance was high, increased exploitation consistently improved subsequent performance. Similar results emerged when using token sort ratio as the measure of exploration.
      With binary measures of exploration (topical transitions or containing the previous prompt), exploitation was associated with higher performance in the next iteration regardless of previous performance bracket. Since these are binary measures, they couldn’t capture the non-linear relationship observed with continuous measures in low-performing groups. Nevertheless, we found that topical transitions (more exploration) led to overall performance decreases, with larger decreases for higher previous performance. Similarly, prompts that included the previous prompt showed higher performance, with larger improvements as previous performance increased.
    • Z-score rescaling: Results were consistent with those for unscaled cosine similarity. The primary difference was that the non-linear relationship between performance and continuous measures of exploitation (TSR ratio and cosine similarity with previous prompt) became more pronounced for the bottom two brackets of previous performance. For low-performing prompts in the previous attempt, moderate exploitation levels yielded optimal performance, with significant performance deterioration at both high and low exploitation levels.
    • Percentile rank rescaling: We found the same general patterns described above, with the non-linearity in exploration/exploitation versus performance in the bottom two brackets of previous performance even stronger than observed with Z-score rescaled cosine similarity.
@@
    • No rescaling: Results were consistent with those described for unscaled cosine similarity.
    • Z-score rescaling: Results matched those described for Z-score cosine similarity, with particularly stark non-linearity between performance and continuous measures of exploitation at the user-prompt level for the bottom two brackets of previous performance.
    • Percentile rank rescaling: Results were consistent with those described for percentile rank rescaled cosine similarity.
These findings suggest a nuanced relationship between exploration/exploitation strategies and performance. Users who were generally more successful tended to employ more exploitative strategies, refining and building upon previous prompts rather than making dramatic changes. However, at the iteration level, the optimal strategy depended on previous performance: when performance was already high, continued exploitation yielded the best results; when performance was poor, moderate exploration was beneficial. This pattern was consistent across different outcome measures and scaling approaches.
@@
Analysis approach: To test this hypothesis, we estimated linear models with each demographic trait as the independent variable, treatment arm fixed effects as controls, and various prompting behaviors (described in Section C.C.5) as the dependent variables. To account for multiple testing, we adjusted p-values across all models using the Benjamini-Hochberg procedure.
Results: Our analysis revealed several significant demographic differences in prompting strategies, with consistent patterns emerging across different measures of prompt similarity and evolution. The detailed findings for each dependent variable are summarized below:
@@
  • Prompt embedding variance: We did not find any significant differences across demographic traits in the overall variance of prompt embeddings, suggesting that the breadth of prompt space exploration was relatively consistent across different demographic groups.
@@
  • Strategic Shifts: We observed statistically significant differences based on age, programming frequency, and instructional writing frequency. Specifically, younger participants, those with low programming frequency, and those with some instructional writing experience demonstrated fewer topical transitions across their prompts, indicating a more consistent approach to prompt development.
@@
  • Successive Prompt Token Sort Ratio: Significant differences emerged by age, education, programming frequency, and imagery/instructional writing frequencies. Older users, those with post-graduate degrees, those with high programming frequency, and those with high writing frequency (both for precise instructions and imagery) wrote prompts that were less similar to their previous attempts as measured by token sort ratio. This suggests these groups took a more exploratory approach to prompt iteration.
@@
  • Successive similarity: We found significant differences across numerous demographic factors: age, gender, education, generative AI outlook, programming skill/frequency, instructional/imagery writing frequency, and certain occupational skills. On average, older users, males, those with post-graduate degrees, frequent programmers, self-reported skilled programmers, those with positive outlooks toward generative AI, frequent generative AI users, those who frequently write instructions or imagery, and those reporting critical thinking and social perceptiveness as occupational skills wrote prompts with lower cosine similarity to their previous attempts. This consistent pattern across multiple demographic factors suggests robust differences in exploration-exploitation tendencies.
@@
  • Successive Prompt ’Contains Previous Prompt’ Dummy: We found significant differences by age, gender, outlook toward generative AI, and imagery writing frequency. Older users and those with neutral outlooks toward generative AI were less likely to write prompts that contained their previous prompt as an exact substring. In contrast, males and those with some imagery writing skills were more likely to build directly upon their previous prompts by including them in subsequent attempts.
These findings reveal a nuanced picture of demographic differences in prompt engineering approaches. Overall, younger users, those with less programming experience, and those with less generative AI experience tended to employ more exploitative strategies, building more consistently upon their previous prompts. In contrast, older users, those with more technical backgrounds, and those with more exposure to generative AI systems exhibited more exploratory behaviors, making larger changes between successive prompts.
Interestingly, while H1 showed that a more exploitative approach was generally associated with better performance, here we find that demographics associated with technical expertise (programming experience, education, etc.) tend toward more exploratory approaches. This apparent contradiction suggests that the relationship between background, prompting strategy, and performance is complex and potentially mediated by other factors not captured in these analyses.
@@
Analysis approach: To test this hypothesis, we examined how the random variation in image quality resulting from model stochasticity affects subsequent user behavior. We first calculated a Z-score for each participant’s generated image relative to the expected distribution for that prompt, following the procedure outlined in Section C.C.3 (with computational details in Section D). This Z-score quantifies how much better or worse the actually generated image was compared to what would be expected for that prompt on average. Higher Z-scores represent instances where the realized image quality exceeded expectations, while lower Z-scores indicate instances where the quality was randomly lower than expected.
To analyze the relationship between this stochasticity and subsequent prompting behavior and performance, we transformed the Z-score into a trichotomous variable with three categories: ”lower-than-expected” (Z-score ¡ -0.45), ”expected” (-0.45 $\leq$ Z-score $\leq$ 0.45), and ”higher-than-expected” (Z-score ¿ 0.45). We then performed two-sample t-tests comparing performance and prompting behavior across these three groups. As a robustness check, we also estimated linear models regressing performance and prompting measures on both the trichotomous Z-score variable and the continuous Z-score, with treatment arm fixed effects. Additionally, we conducted analyses at the user level to determine whether the Z-score of the first prompt affects average performance across all subsequent attempts.
Results: Our analyses revealed several consistent patterns across different outcome measures:
@@
    • No rescaling: We found statistically significant evidence that higher Z-scores for images in previous prompts were associated with increased cosine similarity in subsequent attempts. When comparing performance across the trichotomous Z-score variable, the difference between the top bracket (”higher-than-expected”) and bottom bracket (”lower-than-expected”) was statistically significant, with the top bracket showing better performance in the next attempt. However, we did not find significant differences between the middle bracket and either the top or bottom brackets.
      The linear model did not yield a statistically significant relationship between Z-score and performance in the next iteration, likely due to non-linearity in this relationship across the negative and positive ranges of the Z-score distribution.
      At the user level, we found differences between the top and middle brackets of the first prompt’s Z-score, though these differences were marginally significant (p=0.048) and did not account for multiple testing. Linear models at the user level did not reveal statistically significant relationships between the first prompt’s Z-score and average performance in subsequent attempts.
    • Z-score rescaling: Results were similar to those for unscaled cosine similarity. We found a statistically significant difference between the top and bottom Z-score brackets, with higher performance in the next attempt for the top bracket. No statistically significant effects emerged when comparing the middle bracket with either the top or bottom brackets, or when using linear models. Unlike with the unscaled measure, we found no statistically significant effects at the user level, either when comparing Z-score brackets or when using linear models.
    • Percentile rank rescaling: Results were consistent with those for Z-score rescaled cosine similarity.
@@
    • No rescaling: Results showed similar patterns to CLIP embedding cosine similarity. Higher Z-scores in previous prompts were associated with better performance in subsequent attempts. The difference between the top Z-score bracket and both the bottom and middle brackets was statistically significant, with the top bracket showing higher performance in the next attempt. However, the difference between the bottom and middle brackets was not significant.
      As with cosine similarity, linear models did not yield statistically significant relationships between Z-scores and subsequent performance, likely due to non-linearity. At the user level, we found no statistically significant relationships between the first prompt’s Z-score and performance in subsequent attempts.
    • Z-score rescaling: Results closely matched those for unscaled DreamSim scores, with statistically significant differences between the top bracket and both the bottom and middle brackets, but no significant difference between the bottom and middle brackets. No significant effects emerged from linear models or from user-level analyses.
    • Percentile rank rescaling: Results aligned with those for Z-score rescaled DreamSim, with one exception: at the user-attempt level, we observed a statistically significant difference only between the top and bottom brackets, not between the top and middle brackets.
@@
    • Prompt Length: We found statistically significant evidence that higher Z-scores led to longer subsequent prompts. When using the trichotomous variable, the difference between the top and bottom Z-score brackets was significant, with longer prompts on average in the top bracket. However, we did not observe significant differences between the middle bracket and either the top or bottom brackets. Linear models at the user-attempt level showed a positive and statistically significant relationship, with a one-unit increase in Z-score associated with prompts approximately 0.3 words longer on average. At the user level, the Z-score of the first prompt did not significantly affect the length of subsequent prompts.
    • Successive Similarity: Higher Z-scores were associated with increased similarity between consecutive prompts. The top Z-score bracket showed significantly higher prompt similarity compared to both the bottom and middle brackets, though no significant difference emerged between the bottom and middle brackets. Linear models confirmed a positive and statistically significant relationship between Z-scores and successive prompt similarity. No significant effects were found at the user level.
    • Successive Prompt Token Sort Ratio: Results mirrored those for successive cosine similarity, with higher Z-scores associated with increased token sort ratios between successive prompts. This pattern was consistent across both trichotomous variable comparisons and linear models. No significant effects emerged at the user level.
    • Successive Prompt ’Contains Previous Prompt’ Dummy: Higher Z-scores significantly increased the likelihood that subsequent prompts would contain the previous prompt as an exact substring. The difference between top and bottom Z-score brackets was significant, though differences involving the middle bracket were not. Linear models at the user-attempt level confirmed a positive and statistically significant relationship. No significant effects were found at the user level.
These findings reveal a consistent pattern: when the stochastic nature of generative AI produces unexpectedly high-quality outputs for a given prompt, users tend to build more directly upon that prompt in their subsequent attempts—writing longer prompts that are more similar to and often directly incorporate the previous prompt. This adaptive behavior leads to better performance in subsequent attempts. However, this effect appears to be local rather than global; the quality of the first prompt affects the immediate next attempt but does not significantly influence overall performance or prompting behavior across all subsequent attempts.
This pattern suggests that users are sensitive to and learn from random variation in model outputs, even when they have no way of knowing whether a particularly good or bad result stems from the quality of their prompt or from model stochasticity. The tendency to build upon prompts that produce better-than-expected results represents an intuitive but potentially suboptimal learning strategy, as it may attribute too much importance to random variations rather than focusing on the underlying quality of the prompt itself.
@@
Analysis approach: To test this hypothesis, we conducted pairwise comparisons across the three treatment arms using two-sample t-tests to evaluate differences in both performance outcomes and prompting behaviors. For robustness, we also performed ANOVA tests to examine whether any significant differences existed across all three treatment arms simultaneously. P-values were adjusted for multiple testing using the Benjamini-Hochberg procedure.
Results: Our analyses revealed several key differences in both performance and prompting behavior across the different model conditions:
@@
    • Z-score rescaling: We found statistically significant evidence for performance differences across treatment arms when examining participants’ first attempts, average performance across all attempts, and best attempts (all measured using taskwide Z-scores rather than task-iteration Z-scores).
      For first attempt performance, DALL-E 3 (Verbatim) participants significantly outperformed both DALL-E 3 (Revised) and DALL-E 2 participants, while no statistically significant difference emerged between DALL-E 3 (Revised) and DALL-E 2 participants.
      Average performance across all attempts showed the same pattern: DALL-E 3 (Verbatim) outperformed both other conditions, with no significant difference between DALL-E 3 (Revised) and DALL-E 2.
      When comparing participants’ best attempts, we found statistically significant differences between all three treatment arms in a clear hierarchy: DALL-E 3 (Verbatim) produced better results than DALL-E 3 (Revised), which in turn produced better results than DALL-E 2. ANOVA tests for all three performance measures (first, average, and best attempts) confirmed significant effects of model assignment.
    • Percentile rank rescaling: Results using percentile rank rescaling mirrored those found with Z-score rescaling. DALL-E 3 (Verbatim) participants outperformed both other conditions on first attempt performance and average performance, with no significant differences between DALL-E 3 (Revised) and DALL-E 2 for these metrics. For best attempt performance, all pairwise comparisons revealed significant differences, with DALL-E 3 (Verbatim) outperforming DALL-E 3 (Revised), which outperformed DALL-E 2. ANOVA tests confirmed significant effects of model assignment for all three performance measures.
@@
    • Z-score rescaling: Results using DreamSim with Z-score rescaling showed the same pattern as CLIP embedding cosine similarity. DALL-E 3 (Verbatim) participants significantly outperformed both other conditions on first attempt and average performance, with no significant differences between DALL-E 3 (Revised) and DALL-E 2. For best attempts, we again found the same hierarchical pattern of performance differences across all three conditions. ANOVA tests confirmed significant effects across all performance measures.
    • Percentile rank rescaling: Results with percentile rank rescaling of DreamSim scores were consistent with the Z-score findings. DALL-E 3 (Verbatim) participants outperformed both other conditions on first and average performance, with no differences between DALL-E 3 (Revised) and DALL-E 2. For best attempts, all pairwise comparisons were significant, showing DALL-E 3 (Verbatim) ¿ DALL-E 3 (Revised) ¿ DALL-E 2. ANOVA tests confirmed significant effects across all performance measures.
@@
    • Mean prompt Length: We found statistically significant differences in prompt length across treatment arms. Participants using DALL-E 2 wrote significantly shorter prompts compared to both DALL-E 3 (Revised) and DALL-E 3 (Verbatim) groups. Interestingly, there was no significant difference in prompt length between the two DALL-E 3 variants. ANOVA results confirmed a significant effect of model version on prompt length.
    • Aggregate Similarity: We found no statistically significant differences in the overall variability of prompts across the three treatment arms. ANOVA results confirmed no significant effect of model version on this measure, suggesting that participants explored similar breadths of prompt space regardless of the model they were using.
    • Successive Similarity: Analysis of the average cosine similarity between successive prompts revealed no statistically significant differences across treatment arms. ANOVA results confirmed no significant effect of model version on successive prompt similarity, indicating that the tendency to build upon or deviate from previous prompts was similar across models.
    • Successive Prompt Token Sort Ratio: No statistically significant differences emerged across treatment arms. ANOVA results confirmed no significant effect of model version on this measure of textual similarity between successive prompts.
    • Successive Prompt ”Contains Previous Prompt” Dummy: Examining the probability of a current prompt being a superset of the previous prompt showed no statistically significant differences across treatment arms. ANOVA results confirmed no significant effect of model version on this measure.
These findings paint an interesting picture of how model capacity influences both performance and prompting behavior. In terms of performance, DALL-E 3 (Verbatim) consistently outperformed the other conditions across all metrics and scaling approaches, demonstrating the clear advantage of the more advanced model when used without automated prompt revision. The DALL-E 3 (Revised) condition generally performed worse than DALL-E 3 (Verbatim) but better than DALL-E 2 on best attempt measures, suggesting that automated prompt revision partially degraded the benefits of the more capable model.
For prompting behavior, the primary difference observed was in prompt length: participants using either version of DALL-E 3 wrote significantly longer prompts than those using DALL-E 2. This suggests that users recognized and adapted to the increased capacity of the more advanced model by providing more detailed instructions. However, we found no significant differences across models in measures of prompt similarity or evolution, suggesting that while participants provided more detail when using more capable models, their overall strategies for iterating on prompts remained relatively consistent regardless of model assignment.
@@
Analysis approach: To test this hypothesis, we conducted two types of analyses. First, we performed F-tests comparing the variance of participant performance and prompting behaviors between all three pairs of model conditions (DALL-E 2 vs. DALL-E 3 Revised, DALL-E 3 Revised vs. DALL-E 3 Verbatim, and DALL-E 2 vs. DALL-E 3 Verbatim). Second, we estimated quantile treatment effects (QTEs) between all three pairs of models on participant performance and prompting behaviors. We also visually inspected the QTE patterns to determine whether dispersion/”inequality” was being reduced or increased when participants used different models. For example, positive effects for low quantiles and negative/null effects for high quantiles would indicate inequality reduction.